# No evidence for $H_{0}$ anisotropy from Tully–Fisher or supernova distances

### v1

@@ 4.6 Bulk flow measurement
Figure 8: Magnitude of $\bm{V}_{\rm ext}(\bm{r})$ that is superimposed on the C15 velocity field for a radially varying external flow, compared to the constant $\bm{V}_{\rm ext}$ case. The shaded band indicates the 16th to 84th percentile range. The magnitude rises and then decays towards zero, inconsistent with the linearly growing bulk flow expected from $H_{0}$ anisotropy. For $\bm{V}_{\rm ext}(\bm{r})$, we place spline knots at 0, 20, 40, 60, 800,\,20,\,40,\,60,\,80 and $100~h^{-1}\,\mathrm{Mpc}$, indicated by vertical dashed lines. Above $100~h^{-1}\,\mathrm{Mpc}$ the model uses constant extrapolation.
@@ 5 Discussion and Conclusion
We revisited recent claims of Hubble constant anisotropy from direct distance tracers, made with the CosmicFlows-4 Tully–Fisher relation W1 subsample by B25, and extended the analysis to the 2MTF and SFI++ Tully–Fisher relation samples, as well as to the Pantheon+ and Pantheon+Lane supernova samples (the latter being a reanalysis of Pantheon+ by Lane et al. 2024), all restricted to $z<0.05$. Using a forward-modelling framework that simultaneously fits the distance relation, marginalises over distances, and models peculiar velocities, we tested for a dipole in the zero-point of the Tully–Fisher relation and in the standardised absolute magnitude of supernovae. Since we do not apply an absolute distance calibration, this inferred zero-point represents a degenerate combination of the underlying distance relation zero-point and the Hubble constant.
@@ 5.1 Summary of results
In the CosmicFlows-4 W1 Tully–Fisher relation sample we infer a zero-point dipole amplitude of $0.087\pm 0.019$ mag, with more than $2\sigma$ preference for a non-zero dipole magnitude in SFI++, Pantheon+, and Pantheon+Lane as well. Only in the 2MTF sample, which is also the least constraining, do we not infer a significant zero-point dipole. Under the assumption of the underlying zero-point being isotropic and the dipole arising from variation in $H_{0}$ alone (see Eq. 9), the CosmicFlows-4 result corresponds to anisotropies in $H_{0}$ of approximately 4.1 per cent; see Table˜2 for a summary of the main results. We have confirmed that these results are robust to variations in the Galactic extinction correction. Specifically, we tested both the original E(BV)E(B\!-\!V) values used in CosmicFlows-4 (Schlegel et al., 1998) and the maps of Chiang (2023) or Planck Collaboration et al. (2016), jointly sampling the extinction coefficient $R_{\rm W1}$ in each case.
@@
Such a dipole in $H_{0}$ can equivalently be interpreted as a bulk flow that grows linearly with distance. To test if this is the preferred radial dependence we introduce a non-parametric model of radially varying $\bm{V}_{\rm ext}$ on top of the C15 field. We find that its magnitude increases roughly linearly up to $50~h^{-1}\,\mathrm{Mpc}$ but then rapidly decays towards zero (see Fig.˜8), inconsistent with any $H_{0}$ anisotropy. The inferred bulk flow from this radially varying $\bm{V}_{\rm ext}$ remains consistent with $\Lambda$-cold dark matter at all radii, decaying with $r$ as expected (see Fig.˜7). Thus, while the data favours an anisotropic model over one with a constant $\bm{V}_{\rm ext}$, a non-linearly-rising bulk flow is preferred which does not challenge $\Lambda$-cold dark matter. The preference for a dipole in $H_{0}$ is caused by the relative inflexibility of that model and limited radial range of the data.
Moreover, we caution that beyond the Galactic extinction already considered, other angular selection effects could mimic a dipole-like signal. Since our analysis constrains only a degenerate combination of the dipole in the distance-tracer zero-point and the Hubble constant, the detected dipole could in principle reflect the former rather than a cosmological origin. To mitigate this, we tested several independent datasets with different calibrations and found consistent hints of a dipole in both the Tully–Fisher relation and supernovae. For instance, in the case of the CosmicFlows-4 W1 sample, B25 found no evidence of anisotropy in the linewidths, while the use of consistent WISE photometry argues against spurious calibration dipoles. Nevertheless, because the cosmological principle is a foundational principle of the standard cosmological model, extraordinary evidence is required for a robust detection, warranting further tests with future data.
@@ 5.2 Comparison to previous work
We compare our results for the CosmicFlows-4 W1 Tully–Fisher relation sample with those of B25, who also employ the C15 reconstruction to account for peculiar velocities and report a dipole of magnitude $0.063\pm 0.016$ mag in the direction of $(\ell,b)=(142\pm 30^{\circ},52\pm 10^{\circ})$, in good agreement with our measurement, though slightly lower and with smaller uncertainty. They find only weak support for the anisotropic model, reporting odds of 3:13\!:\!1 in favour of the zero-point dipole. The discrepancy in the dipole magnitude likely stems from a combination of the following factors:
@@
    Velocity assignment. While our model queries the velocity field in real space, assigning velocities based on inferred real-space distances, B25 work in redshift space, first mapping the C15 field into redshift space using the prescription of Carr et al. (2022), and then assigning peculiar velocities as a function of the observed redshift of each galaxy (while approximately undoing the fiducial $\bm{V}_{\rm ext}$ and $\beta^{\star}$ of C15 and applying new ones). The observed redshift together with the peculiar velocity then yields an estimate of the cosmological redshift with a scatter of $\sigma_{v}$. This approach introduces complications in the so-called triple-valued zones (Strauss & Willick, 1995), where a given observed redshift corresponds to multiple possible distances along the line of sight. In such regions, the mapping from redshift to peculiar velocity becomes non-unique, whereas the real-space assignment used in our framework retains a one-to-one correspondence between distance and peculiar velocity. B25 use the observed data to infer latent model parameters (cosmological redshift), which are then used to predict the apparent magnitude. Our method more straightforwardly forward-models the observables and hence avoids these complications.
@@
    Lower redshift cut applied to the W1 sample. While B25 impose $cz_{\rm CMB}>3000~\mathrm{km}\,\mathrm{s}^{-1}$, we do not apply such a cut in our main analysis. We verify that applying the same cut yields a dipole magnitude of $0.062\pm 0.022$ in the direction $(\ell,b)=(137\pm 19^{\circ},24\pm 12^{\circ})$, both of which are in a better agreement with B25 than our fiducial result.
@@
    Photometry-quality cut. B25 do not apply a photometry-quality cut, thus retaining approximately $1300$ more galaxies in the CosmicFlows-4 Tully–Fisher relation W1 sample. We test repeating our fiducial analysis without applying the quality cut and recover a dipole of magnitude of $0.066\pm 0.018$ mag in the direction $(\ell,b)=(128\pm 14^{\circ},18\pm 10^{\circ})$. This magnitude is in good agreement with B25, though the direction remains slightly off.
Thus, after matching the lower redshift or photometry-quality cut, we find a dipole in reasonable agreement with B25, though the direction remains mildly inconsistent. The residual differences with B25 are most likely due to the velocity-assignment scheme.
@@
In our analysis we infer a dipole in the degenerate quantity $a_{\rm TFR}+5\log h$ (or $M_{\rm SN}+5\log h$). Interpreted as a dipole in $H_{0}$, this implies that along the dipole direction $H_{0}$ is enhanced, requiring a negative peculiar velocity to reproduce the same observed redshift. Since the aforementioned studies report a dipole in peculiar velocity, we compare the opposite of their inferred directions with our result. Working in the cosmic microwave background frame, we compare our inferred dipole with literature estimates in Fig.˜4. The dipole direction we find in CosmicFlows-4 agrees with both Migkas et al. (2021) and Watkins et al. (2023), the latter using a larger set of CosmicFlows-4 data. By contrast, the Pantheon+ dipole direction disagrees with our Tully–Fisher relation sample but matches well with the Pantheon+ analysis of Sorrenti et al. (2023b), which may reflect the different redshift ranges probed.
On the supernova side, several studies find significant or semi-significant anisotropies (e.g. Hu et al. 2024; Horstmann et al. 2022; Sorrenti et al. 2023a; Sah et al. 2024; Mc Conville & Ó Colgáin 2023; Krishnan et al. 2022; Zhai & Percival 2022; Rahman et al. 2022; Cowell et al. 2023; Sorrenti et al. 2025). Rahman et al. (2022) investigated a possible dipole in the distance moduli of Type Ia supernovae using the JLA sample (Betoule et al., 2014). They applied the reconstruction of C15 to correct for peculiar velocities and found no evidence for anisotropic expansion. However, other than this study, typically these supernovae studies are at higher redshift than we consider; we have cut our sample to be consistent with the Tully–Fisher relation data and to remain within the volume covered by the C15 reconstruction ($z<0.05$). In addition the methodologies differ: instead of fitting dipoles (or multipoles) some studies perform separate analyses in different parts of the sky, which reduces statistical constraining power but enables investigation of more general types of anisotropy. A limiting case of this is the “hemispherical fitting” method which compares results in one half of the sky to the other. Sometimes the fitting is done in frames other than the cosmic microwave background frame (e.g. Sah et al. 2024), sometimes anisotropy specifically in the direction of the cosmic microwave background dipole is searched for (e.g. Zhai & Percival 2022; Krishnan et al. 2022), and sometimes more sophisticated frameworks are employed (e.g. Kalbouneh et al. 2023; Heinesen 2023). The situation is therefore still unclear as to whether supernovae evince a bona fide anisotropy (though we find evidence for it with the Pantheon+ samples). We caution however that such results are sensitive to the peculiar velocity field, which is not well known beyond $z\approx 0.05$.
@@ 5.3 Disentangling the cosmological dipole from coherent flows
Accounting for peculiar velocities assuming the C15 reconstruction may (partially) absorb a potential cosmological $H_{0}$ dipole, making the disentanglement of a cosmological dipole in $H_{0}$ from coherent flows challenging. In principle, anisotropies in the expansion rate would imprint on the density field inferred from redshift-space positions, thereby biasing the reconstructed peculiar velocities. While the C15 reconstruction assumes linear theory and $\Lambda{\rm CDM}$ to relate the density and velocity fields via the continuity equation and growth rate, it does not constrain the amplitude or anisotropy of the density fluctuations beyond the data itself. Consequently, any non-standard large-scale anisotropy in expansion could propagate into the reconstructed velocity field and be inadvertently subtracted. Nevertheless, the measured clustering amplitude of the 2M++ galaxies remains consistent with $\Lambda{\rm CDM}$ expectations. The present analysis therefore searches for any dipolar excess beyond that predicted by C15 based on the 2M++ galaxy positions.
To conclude, using both Tully–Fisher relation galaxy samples and Type Ia supernova samples, we find evidence for anisotropy in the zero-point calibration of the distance-tracer relation. In the CosmicFlows-4 and Pantheon+ sample, the inferred variation in the Hubble constant is at the level of $4.1\pm 0.9$ and $2.3\pm 0.06$ per cent, respectively. Only in the less constraining Tully–Fisher relation samples (2MTF and SFI++) do we find no preference between the isotropic and anisotropic model. These results are robust to variations in Galactic extinction but are sensitive to the treatment of peculiar velocities, which are non-negligible at $z<0.05$. Without accounting for peculiar velocities, the CosmicFlows-4 dipole would be even larger, while the Pantheon+ remains largely unaffected. We have also tested the reanalysis of Pantheon+ by Lane et al. (2024), which yields a dipole larger by $1\sigma$ compared to the fiducial Pantheon+ data, corresponding to a variation in the Hubble constant of $2.8\pm 0.6$ per cent.
While the preference for the zero-point dipole could in principle indicate a genuine anisotropy in the Hubble constant, a more plausible explanation is that it reflects local peculiar velocities not captured by C15 (or systematics) rather than a true cosmological anisotropy. Allowing $\bm{V}_{\rm ext}$ to vary radially, we find its magnitude rises initially and then rapidly decays to zero, yielding a bulk flow fully consistent with $\Lambda$-cold dark matter.
@@
The Carrick et al. (2015) reconstruction is available at https://cosmicflows.iap.fr. The public CosmicFlows-4 data is available at https://edd.ifa.hawaii.edu/dfirst.php. The code and all other data will be made available on reasonable request to the authors.
@@ Acknowledgements
RS acknowledges financial support from STFC Grant No. ST/X508664/1, the Snell Exhibition of Balliol College, Oxford, and the CCA Pre-doctoral Program. HD is supported by a Royal Society University Research Fellowship (grant no. 211046).
@@ A.1 Tully–Fisher flow model
In this section, we summarise the flow model used in this analysis, which is based on our previous work in S25. For a more detailed description, we refer the reader to that study. We employ a forward model that jointly calibrates the Tully–Fisher relation (and its dipole) while also calibrating the local Universe reconstruction (which accounts for smaller-scale peculiar velocity modes) and other model parameters. For each galaxy, we have observations of its redshift, $z_{\rm CMB}$ (converted to the cosmic microwave background frame) along with its uncertainty, $\sigma_{z}$. We define its value multiplied by the speed of light as $\sigma_{\rm cz}\equiv c\sigma_{\rm z}$. The total redshift of a galaxy in the cosmic microwave background frame is given by
@@
where $\hat{\bm{r}}$ is the galaxy line-of-sight unit vector, and $\bm{V}_{\rm ext}$ accounts for external flows originating beyond the reconstruction volume. The parameter $\beta^{\star}$ is a calibration factor to scale the velocities predicted by C15 and is a function of both cosmology and the galaxy bias. When assuming no underlying velocity field, we effectively set $\bm{v}=0$, modelling the flow solely as $\bm{V}_{\rm ext}$. A third parameter, $\sigma_{v}$, which we introduce later, captures small-scale velocity dispersion not accounted for by the reconstruction.
For each galaxy we observe the apparent magnitude $m_{\rm obs}$ with uncertainty $\sigma_{\rm m}$ and the linewidth $\eta_{\rm obs}$ with uncertainty $\sigma_{\eta}$ (which is a distance-independent observable). These quantities constrain the galaxy distance. In S25 we introduce two latent parameters per galaxy to be inferred: the distance $r$ and the true linewidth $\eta_{\rm true}$. The Tully–Fisher relation relates $\eta_{\rm true}$ to the absolute magnitude $M$ via Eq.˜2. Converting $r$ to a distance modulus $\mu$ yields a predicted apparent magnitude, $m_{\rm pred}=\mu(r)+M(\eta_{\rm true})$, with the distance modulus being
@@
We jointly infer the velocity field calibration parameters, ($\bm{V}_{\rm ext},\beta,\sigma_{v}$), the distance indicator parameters, ($a_{\rm TFR},b_{\rm TFR},c_{\rm TFR},\sigma_{\rm int}$), the mean $\hat{\eta}$ and standard deviation $w_{\eta}$ of the $\eta_{\rm true}$ prior, and potentially the zero-point dipole $\bm{\Delta}_{\rm ZP}$. Assuming independent sources, S25 formulate the likelihood of the observed redshift, magnitude, and linewidth given the distance, true linewidth, and the set of model parameters $\bm{\theta}$ as
@@
where $p(S=1\mid\hat{\eta},\,w_{\eta})$ is the expected fraction of retained sources given a truncation in $\eta_{\rm obs}$. Strictly speaking, $p(S=1\mid\hat{\eta},\,w_{\eta})$ is not part of the data likelihood. It should multiply the product of the per-sample likelihoods and the prior as $\left[p(S=1\mid\hat{\eta},\,w_{\eta})\right]^{-n}$, where $n$ is the number of galaxies in the sample. For notational convenience, however, we absorb it into the likelihood (see Kelly et al. 2008 for more details). The first term in Eq.˜28 is the likelihood of the observed redshift given the predicted value, which depends on the inferred distance and peculiar velocity, while the second term is the likelihood of the observed magnitude given the predicted apparent magnitude. Furthermore, we have that
@@
where $\pi(r\mid\bm{\theta})$ is defined in Eq.˜6. Rather than modelling the full HI selection of the Tully–Fisher relation sample, which is the primary reason why the Tully–Fisher relation samples do not extend to higher redshifts, we follow Lavaux (2016) and adopt an effective treatment by setting the distance prior to Eq.˜6, which incorporates both the homogeneous and inhomogeneous Malmquist bias. We may either sample $r$ and $\eta_{\rm true}$ for each galaxy directly with an Hamiltonian Monte Carlo sampler, or marginalise over them numerically at each Markov Chain Monte Carlo step. The former is computationally faster, while the latter yields a lower-dimensional parameter space suitable for evidence computation. Because evidence values are central to this work, we adopt the latter approach and evaluate the two-dimensional integral numerically on a grid in $r$ and $\eta_{\rm true}$ at each Markov Chain Monte Carlo step.

### v2

@@ 4.6 Bulk flow measurement
Figure 8: Magnitude of $\bm{V}_{\rm ext}(\bm{r})$ that is superimposed on the Carrick_2015 velocity field for a radially varying external flow, compared to the constant $\bm{V}_{\rm ext}$ case. The shaded band indicates the 16th to 84th percentile range. The magnitude rises and then decays towards zero, inconsistent with the linearly growing bulk flow expected from $H_{0}$ anisotropy. For $\bm{V}_{\rm ext}(\bm{r})$, we place spline knots at 0, 20, 40, 60, 800,\,20,\,40,\,60,\,80 and $100~h^{-1}\,\mathrm{Mpc}$, indicated by vertical dashed lines. Above $100~h^{-1}\,\mathrm{Mpc}$ the model uses constant extrapolation.
@@ 5 Discussion and Conclusion
We revisited recent claims of Hubble constant anisotropy from direct distance tracers, made with the CosmicFlows-4 Tully–Fisher relation W1 subsample by Boubel_2025, and extended the analysis to the 2MTF and SFI++ Tully–Fisher relation samples, as well as to the Pantheon+ and Pantheon+Lane supernova samples (the latter being a reanalysis of Pantheon+ by Lane_2024), all restricted to $z<0.05$. Using a forward-modelling framework that simultaneously fits the distance relation, marginalises over distances, and models peculiar velocities, we tested for a dipole in the zero-point of the Tully–Fisher relation and in the standardised absolute magnitude of supernovae. Since we do not apply an absolute distance calibration, this inferred zero-point represents a degenerate combination of the underlying distance relation zero-point and the Hubble constant.
@@ 5.1 Summary of results
In the CosmicFlows-4 W1 Tully–Fisher relation sample we infer a zero-point dipole amplitude of $0.087\pm 0.019$ mag, with more than $2\sigma$ preference for a non-zero dipole magnitude in SFI++, Pantheon+, and Pantheon+Lane as well. Only in the 2MTF sample, which is also the least constraining, do we not infer a significant zero-point dipole. Under the assumption of the underlying zero-point being isotropic and the dipole arising from variation in $H_{0}$ alone (see Eq. 9), the CosmicFlows-4 result corresponds to anisotropies in $H_{0}$ of approximately 4.1 per cent; see Table˜2 for a summary of the main results. We have confirmed that these results are robust to variations in the Galactic extinction correction. Specifically, we tested both the original E(BV)E(B\!-\!V) values used in CosmicFlows-4 (Schlegel_1998) and the maps of Chiang_2023 or Planck_2016, jointly sampling the extinction coefficient $R_{\rm W1}$ in each case.
@@
Such a dipole in $H_{0}$ can equivalently be interpreted as a bulk flow that grows linearly with distance. To test if this is the preferred radial dependence we introduce a non-parametric model of radially varying $\bm{V}_{\rm ext}$ on top of the Carrick_2015 field. We find that its magnitude increases roughly linearly up to $50~h^{-1}\,\mathrm{Mpc}$ but then rapidly decays towards zero (see Fig.˜8), inconsistent with any $H_{0}$ anisotropy. The inferred bulk flow from this radially varying $\bm{V}_{\rm ext}$ remains consistent with $\Lambda$-cold dark matter at all radii, decaying with $r$ as expected (see Fig.˜7). Thus, while the data favours an anisotropic model over one with a constant $\bm{V}_{\rm ext}$, a non-linearly-rising bulk flow is preferred which does not challenge $\Lambda$-cold dark matter. The preference for a dipole in $H_{0}$ is caused by the relative inflexibility of that model and limited radial range of the data.
Moreover, we caution that beyond the Galactic extinction already considered, other angular selection effects could mimic a dipole-like signal. Since our analysis constrains only a degenerate combination of the dipole in the distance-tracer zero-point and the Hubble constant, the detected dipole could in principle reflect the former rather than a cosmological origin. To mitigate this, we tested several independent datasets with different calibrations and found consistent hints of a dipole in both the Tully–Fisher relation and supernovae. For instance, in the case of the CosmicFlows-4 W1 sample, Boubel_2025 found no evidence of anisotropy in the linewidths, while the use of consistent WISE photometry argues against spurious calibration dipoles. Nevertheless, because the cosmological principle is a foundational principle of the standard cosmological model, extraordinary evidence is required for a robust detection, warranting further tests with future data.
@@ 5.2 Comparison to previous work
We compare our results for the CosmicFlows-4 W1 Tully–Fisher relation sample with those of Boubel_2025, who also employ the Carrick_2015 reconstruction to account for peculiar velocities and report a dipole of magnitude $0.063\pm 0.016$ mag in the direction of $(\ell,b)=(142\pm 30^{\circ},52\pm 10^{\circ})$, in good agreement with our measurement, though slightly lower and with smaller uncertainty. They find only weak support for the anisotropic model, reporting odds of 3:13\!:\!1 in favour of the zero-point dipole. The discrepancy in the dipole magnitude likely stems from a combination of the following factors:
@@
    Velocity assignment. While our model queries the velocity field in real space, assigning velocities based on inferred real-space distances, Boubel_2025 work in redshift space, first mapping the Carrick_2015 field into redshift space using the prescription of Carr_2022, and then assigning peculiar velocities as a function of the observed redshift of each galaxy (while approximately undoing the fiducial $\bm{V}_{\rm ext}$ and $\beta^{\star}$ of Carrick_2015 and applying new ones). The observed redshift together with the peculiar velocity then yields an estimate of the cosmological redshift with a scatter of $\sigma_{v}$. This approach introduces complications in the so-called triple-valued zones (StraussWillick_1995), where a given observed redshift corresponds to multiple possible distances along the line of sight. In such regions, the mapping from redshift to peculiar velocity becomes non-unique, whereas the real-space assignment used in our framework retains a one-to-one correspondence between distance and peculiar velocity. Boubel_2025 use the observed data to infer latent model parameters (cosmological redshift), which are then used to predict the apparent magnitude. Our method more straightforwardly forward-models the observables and hence avoids these complications.
@@
    Lower redshift cut applied to the W1 sample. While Boubel_2025 impose $cz_{\rm CMB}>3000~\mathrm{km}\,\mathrm{s}^{-1}$, we do not apply such a cut in our main analysis. We verify that applying the same cut yields a dipole magnitude of $0.062\pm 0.022$ in the direction $(\ell,b)=(137\pm 19^{\circ},24\pm 12^{\circ})$, both of which are in a better agreement with Boubel_2025 than our fiducial result.
@@
    Photometry-quality cut. Boubel_2025 do not apply a photometry-quality cut, thus retaining approximately $1300$ more galaxies in the CosmicFlows-4 Tully–Fisher relation W1 sample. We test repeating our fiducial analysis without applying the quality cut and recover a dipole of magnitude of $0.066\pm 0.018$ mag in the direction $(\ell,b)=(128\pm 14^{\circ},18\pm 10^{\circ})$. This magnitude is in good agreement with Boubel_2025, though the direction remains slightly off.
Thus, after matching the lower redshift or photometry-quality cut, we find a dipole in reasonable agreement with Boubel_2025, though the direction remains mildly inconsistent. The residual differences with Boubel_2025 are most likely due to the velocity-assignment scheme.
@@
In our analysis we infer a dipole in the degenerate quantity $a_{\rm TFR}+5\log h$ (or $M_{\rm SN}+5\log h$). Interpreted as a dipole in $H_{0}$, this implies that along the dipole direction $H_{0}$ is enhanced, requiring a negative peculiar velocity to reproduce the same observed redshift. Since the aforementioned studies report a dipole in peculiar velocity, we compare the opposite of their inferred directions with our result. Working in the cosmic microwave background frame, we compare our inferred dipole with literature estimates in Fig.˜4. The dipole direction we find in CosmicFlows-4 agrees with both Migkas_2021 and Watkins_2023, the latter using a larger set of CosmicFlows-4 data. By contrast, the Pantheon+ dipole direction disagrees with our Tully–Fisher relation sample but matches well with the Pantheon+ analysis of Sorrenti_2023, which may reflect the different redshift ranges probed.
On the supernova side, several studies find significant or semi-significant anisotropies, (e.g. Hu; SN_1; SN_2; Sah; Conville; Krishnan; Zhai; Rahman_2022; Cowell_2023; Sorrenti_2025). Rahman_2022 investigated a possible dipole in the distance moduli of Type Ia supernovae using the JLA sample (Betoule_2014). They applied the reconstruction of Carrick_2015 to correct for peculiar velocities and found no evidence for anisotropic expansion. However, other than this study, typically these supernovae studies are at higher redshift than we consider; we have cut our sample to be consistent with the Tully–Fisher relation data and to remain within the volume covered by the Carrick_2015 reconstruction ($z<0.05$). In addition the methodologies differ: instead of fitting dipoles (or multipoles) some studies perform separate analyses in different parts of the sky, which reduces statistical constraining power but enables investigation of more general types of anisotropy. A limiting case of this is the “hemispherical fitting” method which compares results in one half of the sky to the other. Sometimes the fitting is done in frames other than the cosmic microwave background frame (e.g. Sah), sometimes anisotropy specifically in the direction of the cosmic microwave background dipole is searched for (e.g. Zhai; Krishnan), and sometimes more sophisticated frameworks are employed (e.g. Basheer; Heinesen). The situation is therefore still unclear as to whether supernovae evince a bona fide anisotropy (though we find evidence for it with the Pantheon+ samples). We caution however that such results are sensitive to the peculiar velocity field, which is not well known beyond $z\approx 0.05$.
@@ 5.3 Disentangling the cosmological dipole from coherent flows
Accounting for peculiar velocities assuming the Carrick_2015 reconstruction may (partially) absorb a potential cosmological $H_{0}$ dipole, making the disentanglement of a cosmological dipole in $H_{0}$ from coherent flows challenging. In principle, anisotropies in the expansion rate would imprint on the density field inferred from redshift-space positions, thereby biasing the reconstructed peculiar velocities. While the Carrick_2015 reconstruction assumes linear theory and $\Lambda{\rm CDM}$ to relate the density and velocity fields via the continuity equation and growth rate, it does not constrain the amplitude or anisotropy of the density fluctuations beyond the data itself. Consequently, any non-standard large-scale anisotropy in expansion could propagate into the reconstructed velocity field and be inadvertently subtracted. Nevertheless, the measured clustering amplitude of the 2M++ galaxies remains consistent with $\Lambda{\rm CDM}$ expectations. The present analysis therefore searches for any dipolar excess beyond that predicted by Carrick_2015 based on the 2M++ galaxy positions.
To conclude, using both Tully–Fisher relation galaxy samples and Type Ia supernova samples, we find (apparent) evidence for anisotropy in the zero-point calibration of the distance-tracer relation. In the CosmicFlows-4 and Pantheon+ sample, the implied variation in the Hubble constant is at the level of $4.1\pm 0.9$ and $2.3\pm 0.06$ per cent, respectively. Only in the less constraining Tully–Fisher relation samples (2MTF and SFI++) do we find no preference between the isotropic and anisotropic model. These results are robust to variations in Galactic extinction but are sensitive to the treatment of peculiar velocities, which are non-negligible at $z<0.05$. Without accounting for peculiar velocities, the CosmicFlows-4 dipole would be even larger, while the Pantheon+ remains largely unaffected. We have also tested the reanalysis of Pantheon+ by Lane_2024, which yields a dipole larger by $1\sigma$ compared to the fiducial Pantheon+ data, corresponding to a variation in the Hubble constant of $2.8\pm 0.6$ per cent.
While the preference for the zero-point dipole could in principle indicate a genuine anisotropy in the Hubble constant, a more plausible explanation is that it reflects local peculiar velocities not captured by Carrick_2015 (or systematics) rather than a true cosmological anisotropy. Allowing $\bm{V}_{\rm ext}$ to vary radially, we find its magnitude rises initially and then rapidly decays to zero, yielding a bulk flow fully consistent with $\Lambda$-cold dark matter.
@@
The Carrick_2015 reconstruction is available at https://cosmicflows.iap.fr. The public CosmicFlows-4 data is available at https://edd.ifa.hawaii.edu/dfirst.php. The code underlying the article is available at https://github.com/Richard-Sti/CANDEL and all other data will be made available on reasonable request to the authors.
@@ Acknowledgements
RS acknowledges financial support from STFC Grant No. ST/X508664/1, the Snell Exhibition of Balliol College, Oxford, and the CCA Pre-doctoral Program. HD is supported by a Royal Society University Research Fellowship (grant no. 211046). GL acknowledges support from the Simons Foundation through the Simons Collaboration on “Learning the Universe”.
@@ A.1 Tully–Fisher flow model
In this section, we summarise the flow model used in this analysis, which is based on our previous work in VF_olympics. For a more detailed description, we refer the reader to that study. We employ a forward model that jointly calibrates the Tully–Fisher relation (and its dipole) while also calibrating the local Universe reconstruction (which accounts for smaller-scale peculiar velocity modes) and other model parameters. For each galaxy, we have observations of its redshift, $z_{\rm CMB}$ (converted to the cosmic microwave background frame) along with its uncertainty, $\sigma_{z}$. We define its value multiplied by the speed of light as $\sigma_{\rm cz}\equiv c\sigma_{\rm z}$. The total redshift of a galaxy in the cosmic microwave background frame is given by
@@
where $\hat{\bm{r}}$ is the galaxy line-of-sight unit vector, and $\bm{V}_{\rm ext}$ accounts for external flows originating beyond the reconstruction volume. The parameter $\beta^{\star}$ is a calibration factor to scale the velocities predicted by Carrick_2015 and is a function of both cosmology and the galaxy bias. When assuming no underlying velocity field, we effectively set $\bm{v}=0$, modelling the flow solely as $\bm{V}_{\rm ext}$. A third parameter, $\sigma_{v}$, which we introduce later, captures small-scale velocity dispersion not accounted for by the reconstruction.
For each galaxy we observe the apparent magnitude $m_{\rm obs}$ with uncertainty $\sigma_{\rm m}$ and the linewidth $\eta_{\rm obs}$ with uncertainty $\sigma_{\eta}$ (which is a distance-independent observable). These quantities constrain the galaxy distance. In VF_olympics we introduce two latent parameters per galaxy to be inferred: the distance $r$ and the true linewidth $\eta_{\rm true}$. The Tully–Fisher relation relates $\eta_{\rm true}$ to the absolute magnitude $M$ via Eq.˜2. Converting $r$ to a distance modulus $\mu$ yields a predicted apparent magnitude, $m_{\rm pred}=\mu(r)+M(\eta_{\rm true})$, with the distance modulus being
@@
We jointly infer the velocity field calibration parameters, ($\bm{V}_{\rm ext},\beta,\sigma_{v}$), the distance indicator parameters, ($a_{\rm TFR},b_{\rm TFR},c_{\rm TFR},\sigma_{\rm int}$), the mean $\hat{\eta}$ and standard deviation $w_{\eta}$ of the $\eta_{\rm true}$ prior, and potentially the zero-point dipole $\bm{\Delta}_{\rm ZP}$. Assuming independent sources, VF_olympics formulate the likelihood of the observed redshift, magnitude, and linewidth given the distance, true linewidth, and the set of model parameters $\bm{\theta}$ as
@@
where $p(S=1\mid\hat{\eta},\,w_{\eta})$ is the expected fraction of retained sources given a truncation in $\eta_{\rm obs}$. Strictly speaking, $p(S=1\mid\hat{\eta},\,w_{\eta})$ is not part of the data likelihood. It should multiply the product of the per-sample likelihoods and the prior as $\left[p(S=1\mid\hat{\eta},\,w_{\eta})\right]^{-n}$, where $n$ is the number of galaxies in the sample. For notational convenience, however, we absorb it into the likelihood (see Kelly_2008 for more details). The first term in Eq.˜28 is the likelihood of the observed redshift given the predicted value, which depends on the inferred distance and peculiar velocity, while the second term is the likelihood of the observed magnitude given the predicted apparent magnitude. Furthermore, we have that
@@
where $\pi(r\mid\bm{\theta})$ is defined in Eq.˜6. Rather than modelling the full HI selection of the Tully–Fisher relation sample, which is the primary reason why the Tully–Fisher relation samples do not extend to higher redshifts, we follow Lavaux_Virbius and adopt an effective treatment by setting the distance prior to Eq.˜6, which incorporates both the homogeneous and inhomogeneous Malmquist bias. We may either sample $r$ and $\eta_{\rm true}$ for each galaxy directly with an Hamiltonian Monte Carlo sampler, or marginalise over them numerically at each Markov Chain Monte Carlo step. The former is computationally faster, while the latter yields a lower-dimensional parameter space suitable for evidence computation. Because evidence values are central to this work, we adopt the latter approach and evaluate the two-dimensional integral numerically on a grid in $r$ and $\eta_{\rm true}$ at each Markov Chain Monte Carlo step. We define a fixed radial distance grid ranging from $0.001$ to $201~h^{-1}\,\mathrm{Mpc}$ with a step size of $0.5~h^{-1}\,\mathrm{Mpc}$, which is sufficient given that the Carrick_2015 field is smoothed on scales of $4~h^{-1}\,\mathrm{Mpc}$. For $\eta_{\rm true}$, we adopt an adaptive binning scheme. The Gaussian likelihood term $\mathcal{L}(\eta_{\rm obs}\mid\eta_{\rm true})$ determines the region of $\eta_{\rm true}$ that carries non-negligible probability mass. Accordingly, for each source we define a grid spanning $\eta_{\rm obs}\pm 5\sigma_{\eta}$, discretised into $31$ equally spaced steps, and evaluate the likelihood over a $402\times 31$ grid at each Markov Chain Monte Carlo iteration. We then marginalise over this two-dimensional grid using Simpson’s rule. This computation is performed on GPUs using our JAX-based implementation [^5] https://github.com/jax-ml/jax. To verify that the numerical integration does not introduce bias, we compared the results to a model in which the latent parameters $\eta_{\rm true}$ were explicitly sampled and found the posteriors to be consistent.